OpenStudy (anonymous):

Someone please confirm ? if this is right or not. Problem: Sara needs to take a taxi to get to the movies. The taxi charges $4.00 for the first time, then $2.75 for each mile after that. If the total charge is $20.50,then how far was Sara's taxi ride to the movie? Equation I came up with is 4.00+2.75X=$20.50 and then subtract from there...

OpenStudy (texaschic101):

you got the equation right.. 4 + 2.75x = 20.50 2.75x = 20.50 - 4 2.75x = 16.50 x = 16.50/2.75 x = 6 check.. 4 + 2.75(6) = 20.50 4 + 16.50 = 20.50 20.50 = 20.50 (correct) Her taxi ride to the movies was 6 miles